OpenStudy (anonymous):

Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power

OpenStudy (anonymous):

\[\left\{ \left( x \right)^{m} \right\}^{n}=x ^{m*n}\] \[\left\{ \left( 5 \right)^{\frac{ 3 }{ 4 }} \right\}^{\frac{ 2 }{ 3 }}\]

OpenStudy (welshfella):

first multiply 3/4 by 2/3 according to the first rule that surjit posted

OpenStudy (welshfella):

what do you get for that

OpenStudy (anonymous):

6/12

OpenStudy (anonymous):

hello @welshfella

OpenStudy (welshfella):

right and that simplifies to ?

OpenStudy (welshfella):

hi

OpenStudy (anonymous):

2

OpenStudy (welshfella):

no - thats 1/2